# Evolution of mutating pathogens in networked populations

### v1

@@ Abstract
Epidemic spreading over populations networks has been an important subject of research for several decades, and especially during the Covid-19 pandemic. Most epidemic outbreaks are likely to create multiple mutations during their spreading over the population. In this paper, we study the evolution of a pathogen which can mutate continuously during the epidemic spreading. We consider pathogens whose mutating parameter is the mortality mean-time, and study the evolution of this parameter over the spreading process. We use analytical methods to compute the dynamic equation of the epidemic and the conditions for it to spread. We also use numerical simulations to study the pathogen flow in this case, and to understand the mutation phenomena. We show that the natural selection leads to less violent pathogens becoming predominant in the population. We discuss a wide range of network structures and show how different effects are manifested in each case.
@@ I Introduction
where $<k>$ is the expectation of the number of links of each node in the network. The “transmissibility” can be also described as a general version of the transmission coefficient.
Epidemic thresholds have been studied in a wide range of cases based on different networks structures and epidemic models. For example, Parshani $et\>al.$ [12] used the method of percolation theory and the SIS model spreading on a random graph, to calculate the probability of reinfection. This enables them to compute the $R_{0}$ parameter, that is the average expected infections of each infected node. The $R_{0}$ used to get the critical values of an epidemic, when for an epidemic occurrence we demand $R_{0}>1$.
@@ III Tree graphs
where $P_{\alpha}$ is the probability for an infected node with pathogen $\alpha$ to infect one of its offspring and $d$ is the distance of the specific layer from the tree’s root. Since the $\alpha$ possible values are: $-(d-1)<\alpha<(d-1)$, the mean total number of infected nodes with distance $d$ from the root is:
@@
[…]
@@ IV Random networks
where $<k>$ is the expectation of the number of links of each node in the network. When a significant part of the nodes in the network have already been infected (long range spreading part), some of the neighbors of an infected node are already infected, and the number of susceptible neighbors no longer behaves exactly like the graph degree distribution.
@@ V.2 Social network
VII conclusion
@@
In this paper we studied the phenomenon of continuous mutating pathogens in networked populations. We show that mutating pathogens have a large effect on the way of epidemics spread among populations. We showed that effect over a wide range of networks and specifically we computed the epidemic equations for continuous mutating pathogens on a binary tree graph as well as the basic reproduction number. We also proved that in the case of mutated pathogens that spread in random networks, there is always a positive probability to reach an endemic state, regardless of the parameters of the initial pathogen. Using numerical simulations, we show how the mutated pathogens spread in many kinds of populated networks and the differences between different structures of networks. The numerical simulations also confirmed the calculations from the analytical parts. The results are consistent with the known effect of the less violent variants that become more dominant over time.

### v2

@@ Abstract
Epidemic spreading over populations networks has been an important subject of research for several decades, and especially during the Covid-19 pandemic. Most epidemic outbreaks are likely to create multiple mutations during their spreading over the population. In this paper, we study the evolution of a pathogen which can mutate continuously during the epidemic spreading. We consider pathogens whose mutating parameter is the mortality mean-time, and study the evolution of this parameter over the spreading process. We use analytical methods to compute the dynamic equation of the epidemic and the conditions for it to spread. We also use numerical simulations to study the pathogen flow in this case, and to understand the mutation phenomena. We show that the natural selection leads to less violent pathogens becoming predominant in the population. We discuss a wide range of network structures and show how different effects are manifested in each case. We also applied our theory in the context of the Covid-19 pandemic, using relevant epidemiological data collected for this outbreak. We provided explanations for the variants spreading processes observed throughout this pandemic.
@@ I Introduction
where <k><k> is the expectation of the number of links of each node in the network. The “transmissibility” can be also described as a general version of the transmission coefficient.
Epidemic thresholds have been studied in a wide range of cases based on different networks structures and epidemic models. For example, Parshani etal.et\>al. [12] used the method of percolation theory and the SIS model spreading on a random graph, to calculate the probability of reinfection. This enables them to compute the $R_{0}$ parameter, that is the average expected infections of each infected node. The $R_{0}$ used to get the critical values of an epidemic, when for an epidemic occurrence we demand $R_{0}>1$.
@@ III Tree graphs
where $P_{\alpha}$ is the probability for an infected node with pathogen $\alpha$ to infect one of its offspring and $d$ is the distance of the specific layer from the tree’s root. Since the $\alpha$ possible values are: $(d-1)<\alpha<(d-1)$, the mean total number of infected nodes with distance $d$ from the root is:
@@
$\int_{0}^{t}\frac{1}{\lambda}\exp\left(-\frac{1}{\lambda}\tau\right)\,d\tau\;,$
@@ IV Random networks
where <k><k> is the expectation of the number of links of each node in the network. When a significant part of the nodes in the network have already been infected (long range spreading part), some of the neighbors of an infected node are already infected, and the number of susceptible neighbors no longer behaves exactly like the graph degree distribution.
@@ V.2 Social network
VII Test case – The Covid-19 pandemic
We tested our theory and predictions on the real well-known ‘Corona’ pandemic. This pandemic originated in Wuhan, China in December 2019. The virus responsible for the global break is named SARS-CoV-2 (Severe Acute Respiratory Syndrome Coronavirus 2). Infection with this virus caused the disease known as COVID–19 (Corona Virus Disease 2019). The World Health Organization (WHO) declared The ‘corona’ a global pandemic in March 2020 [40]. From December 2019, as mentioned, the pandemic continued until May 2023 when the WHO ended the Covid–19 global health emergency [41]. Throughout the pandemic, five variants of concern (VOC) of Covid-19 emerged, each of them with its own special characteristics – Alpha, Beta, Gamma, Delta and Omicron. Of the five VOCs, three of them – Alpha, Delta and Omicron – were very infective and transmissible, and shortly after they emerged they reached dominance over other variants that coexisted with them and caused a global pandemic wave. The Alpha variant first emerged in September 2020. It was the first VOC, and appeared after the first strain of Covid-19 variants that originated in Wuhan and referred as to “wild-type” strain. This variant became dominant over the “wild-type” strain, and caused the first world pandemic wave that occurred from December 2020 until early 2021. The Delta variant first emerged in December 2020. This variant reached dominance over the alpha variant and other variants coexisted with it, and caused the second pandemic wave over most of 2021 [42, 43]. The Omicron variant appeared in November 2021. This variant has a very high transmissibility, and shortly after it emerged it took the dominance over the Delta and other variants and caused the third pandemic wave that took place in 2022-2023. The other two VOCs – Beta and Gamma – emerged first in May 2020 and November 2020, respectively. Each of them showed specific properties that made them riskier than variants strains (such as high transmissibility, severe disease, etc.) that existed before they appeared, and for this reason they were designated as VOCs [44, 45].
We applied our theory using real epidemiological measurements of the infection, mortality and recovery mean time of each of the Covid-19 VOCs Alpha, Beta, Delta and Omicron.
The following is a detailed explanation of how we measured the three mean times. In general, to measure these mean times we used some other epidemiological auxiliary measures related to the Covid-19 VOCs, as will be detailed below. These auxiliary measures have been evaluated in many studies and, naturally, their values vary in different studies. Therefore, whenever the evaluation of a measure was different between studies, we consider its value as the mean of the various values that were presented. Here is the way we measured each of three mean times – infection, mortality and recovery:
(i) Measurement of infection mean time: In epidemiological terms, infection mean time is referred to as the measure named generation time – the interval between when an individual becomes infected (primary infection) and when that individual transmits the infection to another (secondary infection). Since there are difficulties in measuring the generation time, it is replaced by the serial interval measure, that is the interval between the appearance of symptoms of the primary infection to the symptoms appearance of the secondary infection [46, 47]. For the Covid-19 VOCs, the serial interval was measured to be as follows: Alpha – 5.2 days [46, 48], Beta – 4.5 days [49], Delta – 3.7 days [46, 49, 50], and Omicron – 3 days [50, 46, 51].
(ii) Measurement of mortality mean time: The mortality mean time which is the mean interval between infection and death, has not been directly studied and evaluated for the Covid-19 VOCs. Therefore, for measuring it we split this interval into two parts that have been studied and measured – the interval between infection and the onset of symptoms known as the incubation period, and the interval between hospitalization due to symptoms and death for individuals who died from the disease. We also considered 1-2 days between the onset of symptoms and hospitalization. For each of the Covid-19 VOCs, the combination of the incubation period and the hospitalization-to-death interval plus 1-2 days between them gives a measure of its mortality mean time. For the Covid-19 VOCs, the incubation period was measured to be as follows: Alpah – 4.5 days [49, 52, 53], Beta – 4.5 days [54, 55, 56], Delta – 4.35 days [53, 50, 52] and Omicron – 3.35 days [57, 58, 50]. The hospitalization-to-death time is: Alpha – 14.3 days, Beta – 4 days, Delta – 12.8 days and Omicron - 16 days [52, 59]. Combining the previous data gives the following measurements for the VOCs mortality mean time: alpha – 20.3 days, Beta – 10 days, Delta – 18.7 days and Omicron – 20.9 days.
(iii) Measurement of recovery mean time: As in the mortality mean time, the recovery mean time, which is the mean interval between infection and recovery, has not been widely studied and directly evaluated. For measuring the recovery mean time, we split this interval, as before, into the incubation period and the time interval from hospitalization due to symptoms to recovery for individuals who recovered from the disease (plus 1-2 days between symptoms onset and hospitalization). The combination of the incubation period and the hospitalization-to-recovery interval plus 1-2 days between them gives a measure of the recovery mean time. For Covid-19 VOCs, the hospitalization-to-recovery time was measured as follows: Alpha – 11 days, Beta – 6 days, Delta - 6 days, Omicron – 3 days [60, 61]. Combining these data with the incubation period detailed above yields measurements of the recovery mean time of Covid-19 VOCs as follows: Alpha – 17 days, Beta – 12 days, Delta – 11.9 days and Omicron – 7.9 days.
In Table 1 we summarize the various measurements for the VOCs as calculated before. Each of the Alpha, Beta, Delta and Omicron variants is represented in a specific raw in the table. The first three columns list the infection mean time ($\lambda$), the mortality mean time ($\gamma$) and the recovery mean time ($r$), respectively. In the fourth column, the infection probability $P_{inf}$ is calculated for each of the variants according to the formula we developed in Eq. (8), dependent on the three mean times $\lambda$, $\gamma$ and $r$. The fifth column is devoted to another important parameter that is the mortality probability $P_{mort}$. This probability refers to an event in which an infected individual dies from the disease. In our model, it occurs when the inter arrival time of the mortality Poisson process occurs before the inter arrival time of the recovery Poisson process. Considering that the inter arrival time of a Poisson process is exponentially distributed and the immemorial of the exponential distribution, we get for the mortality probability as follows: $P_{mort}=\frac{1/\gamma}{1/\gamma+1/r}$. In the fifth column of the table we calculated $P_{mort}$ for each of the VOCs applying this formula.
Table 1: Summary of mean time intervals in the course of infection by variant.
Variant Infection mean time ($\lambda$) Mortality mean time ($\gamma$) Recovery mean time ($r$) $P_{inf}$ $P_{mort}$
Alpha 5.2 20.3 17 0.64 0.46
Beta 4.5 10 12 0.55 0.55
Delta 3.7 18.7 11.9 0.66 0.39
Omicron 3 20.9 7.9 0.66 0.27
An analysis of Table 1 leads to several conclusions regarding the spread of the Corona virus in the world population, as follows:
(i) The infection probability $P_{inf}$ of Alpha – 0.64, Delta – 0.66 and Omicron – 0.66, which are the three variants that caused the three waves of the pandemic, is almost equal. Therefore, we deduce that the replacements of the Alpha wave by the Delta wave and of the Delta wave by the Omicron wave did not result from a creation and emergence of a variant with much higher infection probability than the previous one. On the other hand, the explanation for these replacements is found in the infection mean time column. The infection mean time decreased between variants as the pandemic progressed, from 5.2 days for Alpha to 4.5 days for Delta and finally 3 days for Omicron. The infection mean time of a variant indicates how fast the process of infecting individuals with this variant occurs. It is independent of the infection probability. Accordingly, we deduce that the Delta wave replaced the Alpha wave and the Omicron wave replaced the Delta wave, due to the fact that the rate of Delta infections was much higher than the rate of Alpha infections, and the rate of Omicron infections was much higher than that of Delta infections.
(ii) The mortality probability $P_{mort}$ decreases between the variants Alpha, Delta and Omicron as the pandemic progressed. for Alpha the probability is 0.46, for Delta it is 0.39 and for Omicron it is 0.27. This fact is an indication to the validity of our main argument that the survivability of a variant increases as the variant virulence decreases. In the Covid-19 case, the dominance reached by Delta over Alpha and by Omicron over Delta, is indeed a result of their increase of the infection rate (as detailed above), but here we identify that factor of low mortality probability has also a significant contribution to the survivability and dominance of a variant.
(iii) The variant Beta is different from the other VOCs Alpha, Delta and Omicron. Its infection probability 0.55 is much lower than the infection probability of each of the Alpha, Delta and Omicron which is 0.65 approximately. On the other hand, the mortality probability of the Beta variant that is 0.55, is much greater than this probability for each of the Aplha, Delta and Omicron variants. Therefore, in accordance with our theorem, the low infection probability of the Beta variant explains the facts that the Beta did not cause a wave in the epidemic and could not reach dominance over the variants coexisted with it. On the other hand, its infection probability 0.55 is significant and did allowed the Beta to create strains of individuals that were infected with it. Combining this with the Beta’s second property of a high mortality probability, gives an explanation why this variant is designated as a VOC. Although its probability of infection was relatively low, it was a very virulence variant such that if an individual was infected with the Beta virus, his chance of dying was very high.
VIII conclusion
@@ VII Test case – The Covid-19 pandemic
In this paper we studied the phenomenon of continuous mutating pathogens in networked populations. We show that mutating pathogens have a large effect on the way of epidemics spread among populations. We showed that effect over a wide range of networks and specifically we computed the epidemic equations for continuous mutating pathogens on a binary tree graph as well as the basic reproduction number. We also proved that in the case of mutated pathogens that spread in random networks, there is always a positive probability to reach an endemic state, regardless of the parameters of the initial pathogen. Using numerical simulations, we show how the mutated pathogens spread in many kinds of populated networks and the differences between different structures of networks. The numerical simulations also confirmed the calculations from the analytical parts. The results are consistent with the known effect of the less violent variants that become more dominant over time. In addition, we applied our theory to the case of the Covid-19 pandemic, including analyzing the spreading of the various variants of the primary pathogen over different periods. We provided explanations for the pandemic’s development and the variants that emerged and spread in the population.
@@
Authors contribution
A.I. contributed to the development of the theory, wrote and implemented the simulations and wrote the draft. R.C. contributed to the theory development and reviewed and edited the draft. A.P. contributed to the development of the theory, wrote the ‘Test case – the Covid-19 pandemic’ section, reviewed and edited the draft and the presentation and supervised the project of writing the paper.